OpenStudy (anonymous):

The limit of (sin^3)x/x as x approaches 0

terenzreignz (terenzreignz):

This: \[\Large \lim_{x\rightarrow 0}\frac{\sin^3(x)}{x}\]

OpenStudy (anonymous):

well yeah

terenzreignz (terenzreignz):

Did you know that the limit of a product is the product of the limits... if both limits exist? :D Why don't we make a little adjustment: \[\Large \lim_{x\rightarrow 0}\frac{\sin(x)}{x}\cdot \sin^2(x)\] Does any bit of this look familiar to you?

OpenStudy (anonymous):

\[\frac{ sinx }{x }\]\[=1\]

terenzreignz (terenzreignz):

That's right :) What about the limit of \[\large \sin^2(x)\] as x goes to zero? The sine function is continuous after all...

OpenStudy (anonymous):

So what about \[\sin ^{2}(x)\] ?

terenzreignz (terenzreignz):

It's continuous, so the limit of that as x goes to zero is just \(\large \sin^2(x)\) itself evaluated at x = 0

OpenStudy (anonymous):

alright that makes more sense. Thank you!

terenzreignz (terenzreignz):

So what's your answer?

OpenStudy (anonymous):

Isn't it 0?

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Yeah it's not coming to me

terenzreignz (terenzreignz):

Okay \[\Large \lim_{x\rightarrow 0}\cancel{\color{blue}{\frac{\sin(x)}{x}}}^1\cdot \sin^2(x)\]This bit goes to 1 and... \[\Large \lim_{x\rightarrow 0}\cancel{\color{blue}{\frac{\sin(x)}{x}}}^1\cdot \cancel{\color{red}{\sin^2(x)}}^0\] This bit goes to zero... so all in all?

OpenStudy (anonymous):

it leaves \[\sin ^{2}x\]

terenzreignz (terenzreignz):

Yeah...which goes to...?

OpenStudy (anonymous):

0